# General-Purpose Multicore Architectures

### v1

@@ 4.3 Mitigating Interference
The combination of private caches and shared caches attempts to balance the impact of interference with the need to provision resources for worst-case behavior: private caches allow each CPU core to manage a small working set of its data without first-order impacts of interference, while the shared cache enables resource pooling to reduce the frequency of long-latency misses to main memory for workloads with large data footprints. However, it is still possible for memory interference, at the caches and/or at main memory, to negatively impact one or more cores. In this section, we look at three such examples, as well as potential mechanisms to mitigate them.
@@
Our third example examines how memory scheduling algorithms can introduce unintentional slowdown for threads. Recall from Section 4.2 that the commonly-used FR-FCFS scheduling algorithm does not account for any sort of thread or application awareness, and solely focuses on reducing the number of row activations and precharge operations. While reducing activate and precharge operations help decrease the AMAT for a thread that can exploit row locality, they can unfairly slow down the other threads in a multicore CPU, by continuing to prioritize requests from the thread whose row is already activated. In this example, we have one thread that is generating many loads and stores to the same DRAM row, which is currently activated, while the other threads each have a single load that is waiting to access a different row in the same bank. FR-FCFS will look at all of the queued requests, and prioritize the requests from the first thread because their target row is already activated. This causes all of the requests from the other threads to wait longer in the queue, and in extreme cases can lead to unintentional starvation for these threads. To address this, researchers have proposed a number of memory schedulers that explicitly augment the memory request metadata with a thread ID, and incorporate information about the thread into the scheduling algorithms. As one example, the memory controller can use lightweight runtime metrics to predict which threads are being slowed down due to memory interference, and can prioritize requests from these threads before prioritizing row locality [63].
@@ 4.4 Cache Coherence
In a single-core CPU, write-back caches ensure that the CPU core sees the most recent version of the data for each cache block, as the core always starts memory lookups from the top cache level in the memory hierarchy. However, with multicore CPUs, a dirty cache block stored in one core’s private cache is not immediately visible to other cores in the CPU. This is because if we maintain the same cache lookup procedure as single-core CPUs, a multicore CPU’s core would access its own private caches and the shared last-level cache, and not the private caches of other cores. If left unmodified, this can lead to correctness issues, as one core may ignore and potentially incorrectly overwrite the updates from another core.
@@ 4.5 Memory Consistency Models
A key consideration in coordinating memory updates across cores is the need to present a globally consistent view of these updates to all cores. While cache coherence ensures that updates made by any one core to one unit of data are visible immediately to all cores, memory consistency [65] deals with how we interleave requests across multiple units of data. Unlike cache coherence, which is a purely microarchitectural technique (i.e., transparent to the programmer) because it is a result of only microarchitectural design choices, memory consistency is an architectural technique that must be exposed to the programmer, because data communication across threads is determined by the programmer in both implicit and explicit ways. There are many different memory consistency models, where each memory consistency model defines which possible memory interleavings can be observed by a core. Regardless of the consistency model, each core must maintain an ordering of loads and stores that does not violate true (i.e., read after write) register dependencies, and because caches are part of the shared-memory state, cache coherence is implemented. We examine three popular types of consistency models in this section.
@@
We refer the reader to a detailed discussion [65] for more information about these and other memory consistency models.
@@ 5 Optimizing Operating Systems for Multicore CPUs
As mentioned in Section 2, multicore CPUs can make use of two types of parallelism: multiprocessing of multiple applications, and thread-level parallelism (TLP) within an application. In order to best exploit these types of parallelism on multicore CPU hardware, operating systems and user applications have evolved in a number of key ways. While these changes are not required to make use of multicore CPUs, they can allow developers to significantly improve the overall performance and efficiency of the system. We briefly touch on operating system changes and optimizations for multicore CPUs in this section.
To facilitate our explanations, we will explain these software optimizations from the perspective of threads. As a simple broad definition, a thread is a sequence of CPU instructions from a program, and serves as the basic unit of scheduling for the operating system. Generally speaking, when a program starts executing, that specific executing instance is known as a process, and a process can consist of one or more threads of execution.
@@
Second, as we discuss further in Section 7.2, modern multicore CPUs no longer have homogeneous cores. As a result, the choice of hardware context to assign to a thread can have a significant impact on its performance and energy consumption (e.g., assigning a thread to a big core when it needs only a little core). Early systems with two types of cores made the difference between cores transparent to the OS: a single hardware context was associated with both a big core and a little core, and the hardware would use the CPU frequency setting chosen by the OS for the CPU (see Controlling CPU Core Frequency below) to decide whether the thread should run on the big core or on the little core. Modern systems can have more than two types of CPU cores, and often expose all of the CPU cores directly to the OS. To manage these cores efficiently, OSes now include heterogeneity-aware schedulers, such as the Energy Aware Scheduling approach available in modern versions of the Linux kernel [53].
@@
As we introduced in Section 3.1, multicore CPUs expose the ability to perform per-core dynamic voltage and frequency scaling (DVFS). In modern systems, the voltage and frequency setting for each core is chosen by the OS. CPU manufacturers expose the DVFS capabilities as a series of frequency steps, where the OS chooses a target frequency based on certain properties, and the CPU uses the selected frequency step to control the voltage and frequency of the core.
While specific implementations vary, we will focus on the Linux kernel implementation due to its readily-available documentation. Linux includes a series of governors [89], which are policies that the OS uses to select a frequency setting for a CPU. Some governors, such as performance or powersave, constantly run a CPU at a fixed frequency (for the two examples, maximum or minimum, respectively). Other governors, such as ondemand and conservative, use basic thread statistics such as load and idle time to select the frequency. Newer governors, such as schedutil, extract information about which threads were scheduled for the upcoming scheduling quantum to determine the frequency. In addition to these built-in governors, developers can create custom governor policies for their systems. A recent addition to the Linux kernel, Capacity Aware Scheduling [52], combines core heterogeneity information with frequency settings chosen by the governor to guide hardware context assignment during thread scheduling.
@@
On the application side, there are two ways to take advantage of the parallelism and efficiency offered by multicore CPUs. The first method is to rewrite programs as multithreaded applications. Instead of writing a program as a fully-sequential series of functions, a programmer can identify opportunities to perform some parts of the application concurrently. The programmer can use threading libraries to either (1) explicitly spawn these parallel parts into independent threads, or (2) demarcate regions of the program that inform an advanced library to automatically generate threads. Note that threads do not have to be fully independent, and that programmers can use synchronization primitives to coordinate execution across threads (e.g., locks to protect critical section execution, barriers to synchronize task or sub-task completion), as well as shared memory or message passing to exchange information between threads. While we do not go into multithreaded programming in detail, we refer the reader to other works [58, 22] that provide in-depth coverage of parallel programming techniques and frameworks.
@@ 6 Evaluating Multicore CPUs
While conventional metrics such as speedup have been widely used in the architecture community for decades, several challenges make them difficult to directly apply to multicore CPU evaluations. In this section, we provide a summary of key challenges for performance measurement, and then discuss popular metrics that overcome these challenges. We also briefly discuss metrics related to power and energy, given the emphasis on these throughout the lifetime of multicore CPUs.
@@
During the evaluation of multithreaded applications, an important challenge is factoring in the synchronization overhead. Due to the non-deterministic nature of synchronization, multiple runs of the same program on the same machine, with identical data sets and an identical number of threads, can have different total runtimes, and can have different total instruction counts (e.g., a thread spinning as it waits to acquire a lock will execute additional instructions for each time it checks the lock variable). As a result, it is important to (1) compare equivalent amounts of work performed, as opposed to specific instruction counts (especially in architectural simulators where we tend to execute only parts of a program); and (2) execute the configuration for each data point multiple times, and report a mean (ideally with error bars) instead of the results of a single run.
@@
where $T_{p}$ is the total execution time of the parallel version of the program for $N$ cores, and $T_{s}$ is the total execution time of the sequential version of the program (and not the parallel version with one core). We use the sequential version of the program to ensure that parallel speedup captures the overheads of synchronization. As a result, $S(1)$ can often be less than 1. Note that unlike with single-thread applications, the IPC (instructions per cycle) of a program should not be used as a substitute for execution time, as IPC values can be skewed by synchronization traffic and other non-deterministic behavior.
@@
For multiprogrammed workloads (i.e., a bundle of independent, concurrently executing applications), we are unable to make use of the performance metrics used by multithreaded applications because the different applications are working to complete separate tasks. As a result, the applications in a workload can exhibit significant heterogeneity. For example, we may have a four-application workload, where two of the applications are compute-intensive (and thus have high IPCs), while the other two applications are memory intensive (and thus have low IPCs). If we were to use an aggregate measure of total execution time or the sum of IPCs (e.g., for comparing several potential system improvements to decide which one to implement), we may unfairly bias our improvements to one class of applications (e.g., compute-intensive applications) over the other. To avoid this bias when we measure overall system performance for a multiprogrammed workload, we make use of metrics that attempt to normalize the relative benefits to each application, though there are several ways to perform this normalization [21].
Related to this is the issue of fairness. As the applications in a multiprogrammed workload are independent of each other, one application’s resource usage can generate interference that slows down the other applications that are sharing the multicore CPU, compared to if those applications ran alone on the CPU. For any one application in the workload, we define its slowdown in terms of its performance when running alone on the CPU compared to its performance when the CPU is shared with other applications [45, 63]: [^14] To simplify the discussion of multiprogrammed workload metrics, we assume that each application is single-threaded, which allows us to use IPC in the equations (as there is no synchronization overhead). For multithreaded applications in a multiprogrammed workload, IPC should be replaced by total execution time for that application.
@@
A larger value of $WS$ is better, as it indicates higher system throughput (i.e., lower aggregate impacts of interference). Note that for an $n$-application workload, $WS$ typically ranges between 0 and $n$, indicating that the metric is dependent on the number of applications (and often, by proxy, the number of cores), as this represents a throughput for a specific system. As a result, when comparing two systems and reporting improvements due to a system modification, we typically report a ratio of weighted speedups (which we refer to as WS improvement, although this is sometimes confusingly reported as just WS in several papers):
@@
At a high level, power in a CPU can be broken down into a dynamic component (e.g., the power consumed due to the active switching of transistors to perform work, short-circuit power consumed during gate switching when transistors temporarily connect the high voltage rail to ground due to transistor timing variation) and a static component (e.g., the leakage of power due to imperfections in the switching behavior of a transistor). We briefly discuss components that impact dynamic power in Section 3.1. Note that in the past, dynamic power was orders of magnitude larger than static power, so static power was thought of as a trivial factor in total power consumption. Today, because decades of Dennard scaling translated to significant dynamic power reductions, static power makes up a non-trivial fraction of total CPU power. Power consumption has a direct correlation with thermal dissipation, and is used as a proxy to quantify the heat generated by the CPU. The areal power density, which divides the power by the surface area of the CPU die, is used to determine how aggressive thermal cooling solutions (e.g., heatsinks, liquid cooling, fans) need to be to remove dissipated heat from the CPU and keep the die within safe thermal operating limits. While areal power density is used at design time to provision heat dissipation capacity, the CPU makes use of temperature readings from multiple sensors embedded at various locations in both the CPU chip and the motherboard to dynamically control heat dissipation management, including cooling intensity (e.g., fan speed) and CPU power throttling.
@@ 7 The Evolution of Multicore CPUs
In the years that have elapsed since the introduction of multicore CPUs, there have been a number of innovations to the general microarchitecture described in Section 3. While some of these innovations have been limited to specific manufacturers or models, others have become commonplace across modern CPUs. We highlight three of the most significant shifts in multicore CPU design in this section, and leave the exploration of other innovations as an exercise for the reader. The three shifts that we focus on are finding widespread acceptance in contemporary multicore CPUs: (1) the integration of specialized components on-chip alongside the general-purpose CPU cores into what are known as systems-on-chip (SoCs); (2) the diversification of the constituent cores in a multicore CPU; and (3) the advent of composable chiplets that can allow for the easy integration of many smaller silicon dies in a single chip.
@@ 7.2 Heterogeneous CPU Cores
A seminal study in 2003 illustrated how having a heterogeneous combination of large cores and small cores in a single multicore CPU is significantly more energy efficient than the use of homogeneous (i.e., identical) cores [48]. If all of the cores maintain the same instruction set architecture (ISA), an application can migrate from a large core to a smaller core when it enters a more memory-bound phase, and can migrate back to the larger core when it enters a more compute-bound phase. Arm introduced their first CPU cores designed for a heterogeneous multicore CPU in 2011, under the name big.LITTLE [30]. The first big.LITTLE pairing combined large Cortex-A15 cores with small Cortex-A7 cores, where the Cortex-A15 cores could achieve up to 3.0$\times$ the performance of the Cortex-A7 cores, but the Cortex-A7 could achieve up to 3.8$\times$ the energy efficiency, for selected benchmarks. For what Arm defined as low- to mid-range workloads, the Cortex-A7 was expected to execute these workloads with a dramatic reduction in power consumption, with no expected loss in performance compared to the Cortex-A15. Initially, core heterogeneity was hidden from the system software, with the CPU hardware transparently migrating applications between a big core and a paired little core. This meant that for any big–little pair, one of the cores was idle. Today, core heterogeneity is exposed to the OS scheduler, as we discuss in Section 5.

### v2

@@ 4.3 Mitigating Interference
The combination of private caches and shared caches attempts to balance the impact of interference with the need to provision resources for worst-case behavior: private caches allow each CPU core to manage a small working set of its data without first-order impacts of interference, while the shared cache enables resource pooling to reduce the frequency of long-latency misses to main memory for workloads with large data footprints. However, it is still possible for memory interference, at the caches and/or at main memory, to negatively impact one or more cores. This section looks at three such examples, as well as potential mechanisms to mitigate them.
@@
Our third example examines how memory scheduling algorithms can introduce unintentional slowdown for threads. Recall from Section 4.2 that the commonly-used FR-FCFS scheduling algorithm does not account for any sort of thread or application awareness, and solely focuses on reducing the number of row activations and precharge operations. While reducing activate and precharge operations help decrease the AMAT for a thread that can exploit row locality, they can unfairly slow down the other threads in a multicore CPU, by continuing to prioritize requests from the thread whose row is already activated. In this example, one thread is generating many loads and stores to the same DRAM row, which is currently activated, while the other threads each have a single load that is waiting to access a different row in the same bank. FR-FCFS will look at all of the queued requests, and prioritize the requests from the first thread because their target row is already activated. This causes all of the requests from the other threads to wait longer in the queue, and in extreme cases can lead to unintentional starvation for these threads. To address this, researchers have proposed a number of memory schedulers that explicitly augment the memory request metadata with a thread ID, and incorporate information about the thread into the scheduling algorithms. As one example, the memory controller can use lightweight runtime metrics to predict which threads are being slowed down due to memory interference, and can prioritize requests from these threads before prioritizing row locality [64].
@@ 4.4 Cache Coherence
In a single-core CPU, write-back caches ensure that the CPU core sees the most recent version of the data for each cache block, as the core always starts memory lookups from the top cache level in the memory hierarchy. However, with multicore CPUs, a dirty cache block stored in one core’s private cache is not immediately visible to other cores in the CPU. This is because if the same cache lookup procedure for single-core CPUs is used, a multicore CPU’s core would access its own private caches and the shared last-level cache, and not the private caches of other cores. If left unmodified, this can lead to correctness issues, as one core may ignore and potentially incorrectly overwrite the updates from another core.
@@ 4.5 Memory Consistency Models
A key consideration in coordinating memory updates across cores is the need to present a globally consistent view of these updates to all cores. While cache coherence ensures that updates made by any one core to one unit of data are visible immediately to all cores, memory consistency [66] deals with how requests are interleaved across multiple units of data. Unlike cache coherence, which is a purely microarchitectural technique (i.e., transparent to the programmer) because it is a result of only microarchitectural design choices, memory consistency is an architectural technique that must be exposed to the programmer, because data communication across threads is determined by the programmer in both implicit and explicit ways. There are many different memory consistency models, where each memory consistency model defines which possible memory interleavings can be observed by a core. Regardless of the consistency model, each core must maintain an ordering of loads and stores that does not violate true (i.e., read after write) register dependencies, and because caches are part of the shared-memory state, cache coherence is implemented. This section examines three popular types of consistency models.
@@
Please refer to a detailed discussion [66] for more information about these and other memory consistency models.
@@ 5 Optimizing Operating Systems for Multicore CPUs
As mentioned in Section 2, multicore CPUs can make use of two types of parallelism: multiprocessing of multiple applications, and thread-level parallelism (TLP) within an application. In order to best exploit these types of parallelism on multicore CPU hardware, operating systems and user applications have evolved in a number of key ways. While these changes are not required to make use of multicore CPUs, they can allow developers to significantly improve the overall performance and efficiency of the system. This section briefly touches on operating system changes and optimizations for multicore CPUs.
To facilitate our explanations, these software optimizations will be explained from the perspective of threads. As a simple broad definition, a thread is a sequence of CPU instructions from a program, and serves as the basic unit of scheduling for the operating system. Generally speaking, when a program starts executing, that specific executing instance is known as a process, and a process can consist of one or more threads of execution.
@@
Second, as discussed further in Section 7.2, modern multicore CPUs no longer have homogeneous cores. As a result, the choice of hardware context to assign to a thread can have a significant impact on its performance and energy consumption (e.g., assigning a thread to a big core when it needs only a little core). Early systems with two types of cores made the difference between cores transparent to the OS: a single hardware context was associated with both a big core and a little core, and the hardware would use the CPU frequency setting chosen by the OS for the CPU (see Controlling CPU Core Frequency below) to decide whether the thread should run on the big core or on the little core. Modern systems can have more than two types of CPU cores, and often expose all of the CPU cores directly to the OS. To manage these cores efficiently, OSes now include heterogeneity-aware schedulers, such as the Energy Aware Scheduling approach available in modern versions of the Linux kernel [54].
@@
As introduced in Section 3.1, multicore CPUs expose the ability to perform per-core dynamic voltage and frequency scaling (DVFS). In modern systems, the voltage and frequency setting for each core is chosen by the OS. CPU manufacturers expose the DVFS capabilities as a series of frequency steps, where the OS chooses a target frequency based on certain properties, and the CPU uses the selected frequency step to control the voltage and frequency of the core.
While specific implementations vary, this section will focus on the Linux kernel implementation due to its readily-available documentation. Linux includes a series of governors [90], which are policies that the OS uses to select a frequency setting for a CPU. Some governors, such as performance or powersave, constantly run a CPU at a fixed frequency (for the two examples, maximum or minimum, respectively). Other governors, such as ondemand and conservative, use basic thread statistics such as load and idle time to select the frequency. Newer governors, such as schedutil, extract information about which threads were scheduled for the upcoming scheduling quantum to determine the frequency. In addition to these built-in governors, developers can create custom governor policies for their systems. A recent addition to the Linux kernel, Capacity Aware Scheduling [53], combines core heterogeneity information with frequency settings chosen by the governor to guide hardware context assignment during thread scheduling.
@@
On the application side, there are two ways to take advantage of the parallelism and efficiency offered by multicore CPUs. The first method is to rewrite programs as multithreaded applications. Instead of writing a program as a fully-sequential series of functions, a programmer can identify opportunities to perform some parts of the application concurrently. The programmer can use threading libraries to either (1) explicitly spawn these parallel parts into independent threads, or (2) demarcate regions of the program that inform an advanced library to automatically generate threads. Note that threads do not have to be fully independent, and that programmers can use synchronization primitives to coordinate execution across threads (e.g., locks to protect critical section execution, barriers to synchronize task or sub-task completion), as well as shared memory or message passing to exchange information between threads. While this chapter does not go into multithreaded programming in detail, there are other works [59, 22] that provide in-depth coverage of parallel programming techniques and frameworks.
@@ 6 Evaluating Multicore CPUs
While conventional metrics such as speedup have been widely used in the architecture community for decades, several challenges make them difficult to directly apply to multicore CPU evaluations. This section provides a summary of key challenges for performance measurement, and then discuss popular metrics that overcome these challenges. It also briefly discusses metrics related to power and energy, given their emphasis throughout the lifetime of multicore CPUs.
@@
During the evaluation of multithreaded applications, an important challenge is factoring in the synchronization overhead. Due to the non-deterministic nature of synchronization, multiple runs of the same program on the same machine, with identical data sets and an identical number of threads, can have different total runtimes, and can have different total instruction counts (e.g., a thread spinning as it waits to acquire a lock will execute additional instructions for each time it checks the lock variable). As a result, it is important to (1) compare equivalent amounts of work performed, as opposed to specific instruction counts (especially in architectural simulators that tend to execute only parts of a program); and (2) execute the configuration for each data point multiple times, and report a mean (ideally with error bars) instead of the results of a single run.
@@
where $T_{p}$ is the total execution time of the parallel version of the program for $N$ cores, and $T_{s}$ is the total execution time of the sequential version of the program (and not the parallel version with one core). The equation uses the sequential version of the program to ensure that parallel speedup captures the overheads of synchronization. As a result, $S(1)$ can often be less than 1. Note that unlike with single-thread applications, the IPC (instructions per cycle) of a program should not be used as a substitute for execution time, as IPC values can be skewed by synchronization traffic and other non-deterministic behavior.
@@
Multiprogrammed workloads (i.e., a bundle of independent, concurrently executing applications) are unable to employ the performance metrics used by multithreaded applications because the different applications are working to complete separate tasks. As a result, the applications in a workload can exhibit significant heterogeneity. For example, a four-application workload may have two of the applications that are compute-intensive (and thus have high IPCs), while the other two applications are memory intensive (and thus have low IPCs). If one were to use an aggregate metric of total execution time or the sum of IPCs (e.g., for comparing several potential system improvements to decide which one to implement), the metric may unfairly bias improvements to one class of applications (e.g., compute-intensive applications) over the other. To avoid this bias when measuring overall system performance for a multiprogrammed workload, it is important to use metrics that attempt to normalize the relative benefits to each application, though there are several ways to perform this normalization [21].
Related to this is the issue of fairness. As the applications in a multiprogrammed workload are independent of each other, one application’s resource usage can generate interference that slows down the other applications that are sharing the multicore CPU, compared to if those applications ran alone on the CPU. For any one application in the workload, its slowdown is defined in terms of its performance when running alone on the CPU compared to its performance when the CPU is shared with other applications [46, 64]: [^14] To simplify the discussion of multiprogrammed workload metrics, this section assumes that each application is single-threaded, which allows us to use IPC in the equations (as there is no synchronization overhead). For multithreaded applications in a multiprogrammed workload, IPC should be replaced by total execution time for that application.
@@
A larger value of $WS$ is better, as it indicates higher system throughput (i.e., lower aggregate impacts of interference). Note that for an $n$-application workload, $WS$ typically ranges between 0 and $n$, indicating that the metric is dependent on the number of applications (and often, by proxy, the number of cores), as this represents a throughput for a specific system. As a result, when comparing two systems and reporting improvements due to a system modification, one should typically report a ratio of weighted speedups (occasionally referred to as WS improvement, although this is sometimes confusingly reported as just WS in several papers):
@@
At a high level, power in a CPU can be broken down into a dynamic component (e.g., the power consumed due to the active switching of transistors to perform work, short-circuit power consumed during gate switching when transistors temporarily connect the high voltage rail to ground due to transistor timing variation) and a static component (e.g., the leakage of power due to imperfections in the switching behavior of a transistor). Section 3.1 briefly discusses components that impact dynamic power. Note that in the past, dynamic power was orders of magnitude larger than static power, so static power was thought of as a trivial factor in total power consumption. Today, because decades of Dennard scaling translated to significant dynamic power reductions, static power makes up a non-trivial fraction of total CPU power. Power consumption has a direct correlation with thermal dissipation, and is used as a proxy to quantify the heat generated by the CPU. The areal power density, which divides the power by the surface area of the CPU die, is used to determine how aggressive thermal cooling solutions (e.g., heatsinks, liquid cooling, fans) need to be to remove dissipated heat from the CPU and keep the die within safe thermal operating limits. While areal power density is used at design time to provision heat dissipation capacity, the CPU makes use of temperature readings from multiple sensors embedded at various locations in both the CPU chip and the motherboard to dynamically control heat dissipation management, including cooling intensity (e.g., fan speed) and CPU power throttling.
@@ 7 The Evolution of Multicore CPUs
In the years that have elapsed since the introduction of multicore CPUs, there have been a number of innovations to the general microarchitecture described in Section 3. While some of these innovations have been limited to specific manufacturers or models, others have become commonplace across modern CPUs. This section highlights three of the most significant shifts in multicore CPU design, and leaves the exploration of other innovations as an exercise for the reader. These three shifts are finding widespread acceptance in contemporary multicore CPUs: (1) the integration of specialized components on-chip alongside the general-purpose CPU cores into what are known as systems-on-chip (SoCs); (2) the diversification of the constituent cores in a multicore CPU; and (3) the advent of composable chiplets that can allow for the easy integration of many smaller silicon dies in a single chip.
@@ 7.2 Heterogeneous CPU Cores
A seminal study in 2003 illustrated how having a heterogeneous combination of large cores and small cores in a single multicore CPU is significantly more energy efficient than the use of homogeneous (i.e., identical) cores [49]. If all of the cores maintain the same instruction set architecture (ISA), an application can migrate from a large core to a smaller core when it enters a more memory-bound phase, and can migrate back to the larger core when it enters a more compute-bound phase. Arm introduced their first CPU cores designed for a heterogeneous multicore CPU in 2011, under the name big.LITTLE [31]. The first big.LITTLE pairing combined large Cortex-A15 cores with small Cortex-A7 cores, where the Cortex-A15 cores could achieve up to 3.0$\times$ the performance of the Cortex-A7 cores, but the Cortex-A7 could achieve up to 3.8$\times$ the energy efficiency, for selected benchmarks. For what Arm defined as low- to mid-range workloads, the Cortex-A7 was expected to execute these workloads with a dramatic reduction in power consumption, with no expected loss in performance compared to the Cortex-A15. Initially, core heterogeneity was hidden from the system software, with the CPU hardware transparently migrating applications between a big core and a paired little core. This meant that for any big–little pair, one of the cores was idle. Today, core heterogeneity is exposed to the OS scheduler, as discussed in Section 5.